# artificial intelligence AoA angle-of-arrival

### v1

@@ IV-A Factor Graph Construction and Message Passing Schedule
First, we construct a factor graph for the 3D positioning problem. By introducing an auxiliary random variable $\Gamma_{i}=\mathbf{h}_{i}^{\mathrm{T}}\bm{X}$ for $i=1,\ldots,M$, we write the measurement model (2) as $y_{i}=\gamma_{i}+b_{i}+n_{i}$, where $\gamma_{i}$ is a realization of $\Gamma_{i}$. We form the sets of random variables $\{Y_{i}\}$, $\{B_{i}\}$, $\{\Gamma_{i}\}$, and $\{\Lambda_{i}\}$ into random vectors $\bm{Y}$, $\bm{B}$, $\bm{\Gamma}$, and $\bm{\Lambda}$, and their realizations into vectors $\mathbf{y}$, $\mathbf{b}$, $\bm{\gamma}$, and $\bm{\lambda}$. Following the assumptions in Section II-A, we factorize the joint posterior probability of $\bm{\Gamma}$, $\bm{X}$, $\bm{B}$, and $\bm{\Lambda}$ as shown in equation ()($\ast$)( ∗ ) in Fig. 2, and form a cycle-free factor graph with $M$ branches to represent the factorization, shown in the same figure. Each term in equation ()($\ast$)( ∗ ) corresponds to a factor (function) node in the graph, and for ease of description, we define
@@
In equation ()($\ast$)( ∗ ), $p_{\mathbf{X}}(\mathbf{x})$ represents the prior PDF of $\mathbf{X}$. Since we are focusing on a single epoch, we will omit $p_{\mathbf{X}}(\mathbf{x})$ from message passing.
The sum-product rule [29] is adopted for running the message passing on the cycle-free factor graph, and the schedule is straightforward as follows: Messages are passed from variable node $\Lambda_{i}$ to variable node $\mathbf{X}$ in steps 1- 4 along all $M$ branches. In step 5, variable node $\mathbf{X}$ computes the product of all received messages, which after normalization is the posterior PDF of $\mathbf{X}$. These steps are sufficient for position estimation and PL computation of the current positioning epoch. Optionally, messages can be further passed back to variable node $\Lambda_{i}$ in steps 6- 9 along each branch to obtain the posterior PMF of $\Lambda_{i}$. [^4] In a real-life positioning system, measurement states in consecutive epochs are likely correlated, making it reasonable to update $p_{\Lambda_{i}}(\lambda_{i})$ for the next positioning epoch using the obtained posterior PMF. However, as this is beyond the scope of this paper, so steps 6- 9 are not discussed further.
Figure 2: The factor graph used by the Bayesian RAIM algorithm, constructed according to the joint posterior probability factorization given by ()($\ast$)( ∗ ). The order of message computation and passing is given by the numbers in circles (shown only on the $i$th branch, but is the same for all branches), while the arrows indicate the passing direction.
IV-B General Gaussian Message Computation Rules
@@
For $i=1,\ldots,K$, given a linear mapping $\bm{Y}_{i}=\mathsf{A}_{i}\bm{X}$, where $\bm{X}\in\mathbb{R}^{n}$ and $\bm{Y}_{i}\in\mathbb{R}^{m_{i}}$ are Gaussian random vectors, each matrix $\mathsf{A}_{i}\in\mathbb{R}^{m_{i}\times n}$ is full-rank [^5] This full-rank assumption does not lose generality. If $\mathsf{A}$ is rank-deficient, i.e., $k=\mathrm{rank}(\mathsf{A})<\min(n,m)$, a (rank) decomposition $\mathsf{A}=\mathsf{BC}$ can be obtained (e.g., using compact singular value decomposition (SVD)) where $\mathsf{B}\in\mathbb{R}^{m\times k}$ and $\mathsf{C}\in\mathbb{R}^{k\times n}$ have full rank ($=k$). Thus, $\bm{Y}=\mathsf{A}\bm{X}=\mathsf{B}(\mathsf{C}\bm{X})$ can be treated as two consecutive linear mapping via full-rank matrices., with $\mathrm{rank}(\mathsf{A}_{i})=m_{i}\leq n$. The objective is to infer the PDF of $\bm{X}$ from the known non-degenerate PDF of $\bm{Y}_{i}$ for all $i$, and then compute the product of all inferred PDFs of $\bm{X}$.
By letting $\mathsf{A}_{i}=\mathbf{h}_{i}^{\mathrm{T}}$, our problem becomes a special case of this general problem. A general Gaussian distribution for a random vector $\bm{X}\in\mathbb{R}^{n}$ is characterized by a symmetric positive semi-definite covariance matrix $\mathsf{\Sigma}\succeq 0$ and a mean vector $\mathbf{m}\in\mathbb{R}^{n}$, or equivalently, by $\mathsf{V}\triangleq\mathsf{\Sigma}^{+}$ (the pseudo-inverse of $\mathsf{\Sigma}$) and $\mathbf{u}\triangleq\mathsf{V}\mathbf{m}$. Notably, $\mathrm{rank}(\mathsf{\Sigma})=\mathrm{rank}(\mathsf{V})\leq n$, with the $<$ relation occurring when the distribution is degenerate. The PDF expressed using $(\mathbf{m},\mathsf{\Sigma})$ is denoted $f_{\bm{X}}({\mathbf{x}};\mathbf{m},\mathsf{\Sigma})$ (see (43)), and when using $(\mathbf{u},\mathsf{V})$, it is denoted $f^{\mathrm{E}}_{\bm{X}}({\mathbf{x}};\mathbf{u},\mathsf{V})$ (see (45)). Moreover,
@@
For $m=n$, $\mathsf{V}_{\bm{Y}}\succ 0$ ensures $\mathsf{V}_{\bm{X}}\succ 0$ (since $\mathsf{A}$ is full-rank), making the inverse mapping of $\bm{Y}=\mathsf{A}\bm{X}$ unique and straightforward to compute. In this case, (17a)-(17c) can be directly obtained following [30, Eq. (356), (357)].

### v2

@@ III-A Factor Graph Construction and Message Passing Schedule
First, we introduce an auxiliary random variable $\Gamma_{i}=\mathbf{h}_{i}^{\mathrm{T}}\bm{X}$ for $i=1,\ldots,M$, so that we can rewrite the measurement model (2) as $y_{i}=\gamma_{i}+b_{i}+n_{i}$, where $\gamma_{i}$ is a realization of $\Gamma_{i}$. We form the sets of random variables $\{Y_{i}\}$, $\{B_{i}\}$, $\{\Gamma_{i}\}$, and $\{\Lambda_{i}\}$ into random vectors $\bm{Y}$, $\bm{B}$, $\bm{\Gamma}$, and $\bm{\Lambda}$, and their realizations into vectors $\mathbf{y}$, $\mathbf{b}$, $\bm{\gamma}$, and $\bm{\lambda}$. As shown in Fig. 3, we factorize the joint posterior probability of $\bm{\Gamma}$, $\bm{X}$, $\bm{B}$, and $\bm{\Lambda}$ in equation $(\ast)$ following the assumptions in Section II-A, and form a cycle-free factor graph with $M$ branches to represent the factorization. Each term in equation $(\ast)$ corresponds to a factor (function) node in the graph. For ease of description, we define
@@
In equation $(\ast)$, $p_{\mathbf{X}}(\mathbf{x})$ represents the prior PDF of $\mathbf{X}$. Since we are focusing on a single epoch, we will omit $p_{\mathbf{X}}(\mathbf{x})$ from message passing.
Applying the sum-product rule [31] on this cycle-free factor graph leads to a straightforward message passing schedule: Messages are passed from variable node $\Lambda_{i}$ to variable node $\mathbf{X}$ in steps 1- 4 along all $M$ branches in parallel. Then in step 5, variable node $\mathbf{X}$ computes the product of all received messages, which after normalization is the posterior PDF of $\mathbf{X}$. These steps are sufficient for position estimation and PL computation of the current positioning epoch. Messages can be further passed back to variable node $\Lambda_{i}$ in steps 6- 9 along each branch to obtain the posterior PMF of $\Lambda_{i}$. [^2] In a real-life positioning system, measurement states in consecutive epochs are likely correlated, making it reasonable to update $p_{\Lambda_{i}}(\lambda_{i})$ for the next positioning epoch using the obtained posterior PMF. However, as this is beyond the scope of this paper, so steps 6- 9 are not discussed further.
Figure 3: The factor graph used by the Bayesian RAIM algorithm, constructed according to the joint posterior probability factorization given by $(\ast)$. The order of message computation and passing is given by the numbers in circles (shown only on the $i$th branch, but is the same for all branches), while the arrows indicate the passing direction.
III-B General Gaussian Message Computation Rules
@@
For $i=1,\ldots,K$, given a linear mapping $\bm{Y}_{i}=\mathsf{A}_{i}\bm{X}$, where $\bm{X}\in\mathbb{R}^{n}$ and $\bm{Y}_{i}\in\mathbb{R}^{m_{i}}$ are Gaussian random vectors, each matrix $\mathsf{A}_{i}\in\mathbb{R}^{m_{i}\times n}$ is full-rank [^3] This full-rank assumption does not lose generality. If $\mathsf{A}$ is rank-deficient, i.e., $k=\mathrm{rank}(\mathsf{A})<\min(n,m)$, a (rank) decomposition $\mathsf{A}=\mathsf{BC}$ can be obtained (e.g., using compact singular value decomposition (SVD)) where $\mathsf{B}\in\mathbb{R}^{m\times k}$ and $\mathsf{C}\in\mathbb{R}^{k\times n}$ have full rank (=k=k). Thus, $\bm{Y}=\mathsf{A}\bm{X}=\mathsf{B}(\mathsf{C}\bm{X})$ can be treated as two consecutive linear mapping via full-rank matrices., with $\mathrm{rank}(\mathsf{A}_{i})=m_{i}\leq n$. The objective is to infer the PDF of $\bm{X}$ from the known non-degenerate PDF of $\bm{Y}_{i}$ for all $i$, and then compute the product of all inferred PDFs of $\bm{X}$.
By letting $\mathsf{A}_{i}=\mathbf{h}_{i}^{\mathrm{T}}$, our problem becomes a special case of this general problem. A general Gaussian distribution for a random vector $\bm{X}\in\mathbb{R}^{n}$ is characterized by a symmetric positive semi-definite covariance matrix $\mathsf{\Sigma}\succeq 0$ and a mean vector $\mathbf{m}\in\mathbb{R}^{n}$, or equivalently, by $\mathsf{V}\triangleq\mathsf{\Sigma}^{+}$ (the pseudo-inverse of $\mathsf{\Sigma}$) and $\mathbf{u}\triangleq\mathsf{V}\mathbf{m}$. Notably, $\mathrm{rank}(\mathsf{\Sigma})=\mathrm{rank}(\mathsf{V})\leq n$, with the << relation occurring when the distribution is degenerate. The PDF expressed using $(\mathbf{m},\mathsf{\Sigma})$ is denoted $f_{\bm{X}}({\mathbf{x}};\mathbf{m},\mathsf{\Sigma})$ (see (35)), and when using $(\mathbf{u},\mathsf{V})$, it is denoted $f^{\mathrm{E}}_{\bm{X}}({\mathbf{x}};\mathbf{u},\mathsf{V})$ (see (37)). Moreover,
@@
For $m=n$, $\mathsf{V}_{\bm{Y}}\succ 0$ ensures $\mathsf{V}_{\bm{X}}\succ 0$ (since $\mathsf{A}$ is full-rank), making the inverse mapping of $\bm{Y}=\mathsf{A}\bm{X}$ unique and straightforward to compute. In this case, (9a)-(9c) can be directly obtained following [32, Eq. (356), (357)].